I need help with a problem.
How many cubic yards of concrete are needed to fill a hole 18ft long, 4ft wide, and 9in. thick?
Please help.
Kathy Price

Answer by user_dude2008(1861)

9 in = 3/4 ft

V = 18*4*(3/4) = 54 cubic feet

54/27 = 2 cubic yards


Ans: volume is 2 cubic yards